OpenStudy (anonymous):

Find the slope and y-intercept of the line in the given equation. y=5/6x-14

OpenStudy (mysesshou):

Ah, ok. jumped to here.

OpenStudy (mysesshou):

y=(5/6)*x-14 the format for slope intercept is y=m*x+b which numbers do you think are m and b ?

OpenStudy (anonymous):

i think 14 is b idk what m would be though? hmmmm

OpenStudy (mysesshou):

y=(5/6)*x+(-14) try this

OpenStudy (mysesshou):

what is multiplied by the x just like the m does

OpenStudy (anonymous):

the 5/6?

OpenStudy (mysesshou):

yes recheck your b

OpenStudy (anonymous):

so b is 5/6 does that make m 14

OpenStudy (mysesshou):

y=(5/6)*x+(-14) y=(m)*x+(b) b is negative 14

OpenStudy (mysesshou):

so, b=-14 m=5/6 for y=5/6x-14

OpenStudy (mysesshou):

m = slope b = y intercept

OpenStudy (mysesshou):

makes sense?

OpenStudy (anonymous):

a little lol

OpenStudy (mysesshou):

that part y=5/6x-14, the -14, that is the intercept, because the format of the equation is y=m*x+b, so you need to change the equation to a +b format giving y= m*x+(-___) so that inside ( ) is the b

OpenStudy (anonymous):

so the numbers inside the () is b?

OpenStudy (mysesshou):

yes, that is how we get -14

OpenStudy (mysesshou):

since the format of the equation is + b .. so change your problem to fit the format + something... but yours is minus14, so the minus makes the number negative.

OpenStudy (mysesshou):

like if you took y=(5/6)*x+(-14) and then you can change it back to y=(5/6)*x-14 they are equivalent

OpenStudy (mysesshou):

any better?

OpenStudy (mysesshou):

not sure how else to help

OpenStudy (anonymous):

okay i follow i had to write it out on paper lol

OpenStudy (mysesshou):

that is much simpler than typing. it is also easier for me to write than type this out.. so it is an interesting experience :)

OpenStudy (anonymous):

yes it is

OpenStudy (mysesshou):

and i'm still learning site since i just made my acct a day or two ago.. :) so learning to write the equations in code is nifty too :)

OpenStudy (mysesshou):

so you have two left?

OpenStudy (anonymous):

yes

OpenStudy (mysesshou):

ok. type away